OpenStudy (anonymous):

Find the simple interest . $2,525 deposited at an interest rate of 2.5& for 4 years

OpenStudy (anonymous):

* 2.5 %

OpenStudy (kropot72):

Interest for one year = 2525 * 0.025 = $63.125 For 4 years $63.125 * 4 = $252.50

OpenStudy (anonymous):

Thank You (: !

OpenStudy (kropot72):

You're very welcome :)